OpenStudy (anonymous):

Find a power series representation for f(x)=ln((1+x)/(1−x)).

OpenStudy (anonymous):

The formula for ln(1+x)= (-1)^n-1* x^n/n The form. for ln(1-x)= x?? I know I need to combine them. Just not sure how.

OpenStudy (experimentx):

just add (1-x) at the bottom, if you expand (1-x) binomially, you will end up having another series at each term.

OpenStudy (anonymous):

i would start with the fact that you have \[\ln(1+x)-\ln(1-x)\]

OpenStudy (experimentx):

Oo ... didn't see that ... both was inside log.

OpenStudy (anonymous):

ln(1-x)=\[\sum_{n=0}^{\infty} (-1)^{n-1}\frac{x^n}{n}\]

OpenStudy (anonymous):

Sorry, thats ln(1+x)

OpenStudy (anonymous):

Ooooh, you're saying with that subtraction sign, i can distribute and it'll be posi. and be in the same form?!?

OpenStudy (experimentx):

n = 1

OpenStudy (anonymous):

My denominator is 1?

OpenStudy (experimentx):

start summation from n=1

OpenStudy (anonymous):

The issue is that I have to make a formula first to put in the summation.

OpenStudy (experimentx):

\[ ln(1-x) = \sum_{n=1}^{\infty} \frac{-x^n}{n} \] \[ ln(1+x) - ln(1-x) = \sum_{n=1}^{\infty} \frac{(-1)^n x^n}{n} - \frac{-x^n}{n} \]

OpenStudy (experimentx):

\[ \sum_{n=2i+1, i=0}^{\infty} \frac{-2x^n}{n} \]

OpenStudy (anonymous):

Im not allowed to start the sum at 1, lol.

OpenStudy (experimentx):

Oo.. there was an error, above make that - in summation to + at the middle, I am sure if you can start at 0, since 1/n = undefined at n=0

OpenStudy (anonymous):

@sam30317 you got this?

OpenStudy (anonymous):

Nooo, = ( I got it wrong & still a bit lost. I just ned to know how to fuse them

OpenStudy (anonymous):

ok first off well know series for \(\ln(1+x)\) at x = 0 is \[x-\frac{x^2}{2}+\frac{x^3}{3}-...=\sum (-1)^{k+1}x^k\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now to find series for \(\ln(1-x)\) replace \(x\) by \(-x\) in the above series. the even powers (which are all negative above) obviously can't tell the diffreence between \(x\) and \(-x\) so they stay the same, and the odd powers (which are all positive above) become negative

OpenStudy (anonymous):

in other words is t is identical to the one above except that ALL the terms are negative, \[-x-\frac{x^2}{2}-\frac{x^3}{3}-...=-\sum\frac{x^k}{k}\]

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

now subtract the second from the first, which means of course get rid of the minus sign and add the terms

OpenStudy (anonymous):

the even powers will add up to zero

OpenStudy (anonymous):

and you will get double all the odd ones

OpenStudy (anonymous):

\[2x+\frac{2x^3}{3}+\frac{2x^5}{5}+...\]

OpenStudy (anonymous):

yay just was the wolf got it! so we know it is right

OpenStudy (anonymous):

∑(−1)^k+1 x^k - (x^k/k) ?? I didnt have to solve anything. I just needed to find the formula representation of it. Secondly, Wolf isnt showing what I was looking for. Where do you giys see it?

OpenStudy (anonymous):

the formula representation would be \[2\sum\frac{x^{2k-1}}{2k-1}\]

OpenStudy (anonymous):

>____> Thanks everyone! Ihave to review what you giys told me b/c im a bit brain dead. I have to relearn EVERYTHING taught by Monday.

OpenStudy (anonymous):

good luck! stay up late and drink lots of coffee